# Maximal GW amplitude from bubble collisions in supercooled phase transitions

### v1

@@ V Discussion and conclusion
We find that the spectral shape, amplitude, and peak frequency of the GW signal are largely insensitive to details of the bubble nucleation history once the parameter dependence is factorized as $\Omega_{\rm GW}\propto(\tilde{\beta}/\tilde{H}_{*})^{-2}(\kappa_{*}\alpha_{*})^{2}$, where the subscript * denotes quantities evaluated at the end of the phase transition. In the radiation-dominated scenarios considered here, the ratio satisfies $\tilde{\beta}/\tilde{H}_{*}\!\geq\!3$, implying an upper bound on the GW amplitude even for strongly supercooled transitions. Using $\Delta=\mathcal{O}(10^{-2}\text{--}10^{-1})$ from our results, we obtain a present-day peak bound $\Omega_{\rm GW}h^{2}\lesssim 10^{-7}\,\kappa_{*}^{2}\alpha_{*}^{2}$.
@@
Our analytic expressions also permit controlled expansions—large/small $k$ and the flat-spacetime limit—by expanding the integrands in the relevant small parameters. These analyses, together with the full derivation for the full analytic formula, are presented in a companion paper [56].

### v2

@@ V Discussion and conclusion
We present an analytic formulation of bubble collisions for runaway walls under the thin-wall and envelope approximations, taking into account the effects of cosmic expansion while neglecting the backreaction on the background metric. We find that the spectral shape, amplitude, and peak frequency of the GW signal are largely insensitive to details of the bubble nucleation history once the parameter dependence is factorized as $\Omega_{\rm GW}\propto(\tilde{\beta}/\tilde{H}_{*})^{-2}(\kappa_{*}\alpha_{*})^{2}$, where the subscript * denotes quantities evaluated at the end of the phase transition. In the radiation-dominated scenarios considered here, the ratio satisfies $\tilde{\beta}/\tilde{H}_{*}\!\geq\!3$, implying an upper bound on the GW amplitude even for strongly supercooled transitions. Using $\Delta=\mathcal{O}(10^{-2}\text{--}10^{-1})$ from our results, we obtain a present-day peak bound $\Omega_{\rm GW}h^{2}\lesssim 10^{-7}\,\kappa_{*}^{2}\alpha_{*}^{2}$.
@@
Our analytic expressions also permit controlled expansions, such as large/small $k$ and the Minkowski limit, by expanding the integrands in the relevant small parameters. These analyses, together with the full derivation for the full analytic formula, are presented in a companion paper [60].